# SVIP: Towards Verifiable Inference of Open-source Large Language Models

### v1

@@ D.5 Experimental Details of Adapter Attack
D.6 Experimental Details of Secret Recovery Attack
@@
D.7 Case Study: The Vulnerability of the Simple Protocol Without Secret Mechanism
In this case study, we implement the simple protocol and examine its vulnerability to the direct vector optimization attack described in Section 4.2. We use the SoW representation as the self-labeling function. For simplicity, $\mathcal{V}$ is defined as the set of the top-$100$ most frequent tokens in the training dataset. We use Llama-2-13B as the specified model. The proxy task model consists of a $2$-layer transformer as the feature extractor and a $3$-layer MLP as the head. The model is trained for $8$ epochs with a batch size of $512$ following Eq. (1).
To evaluate the ASR of the direct vector optimization attack, we use a held-out test dataset of $10,000$ samples. Each attack vector $\tilde{z}$ is randomly initialized and optimized over $100$ steps using the Adam optimizer (Kingma, 2014) based on Eq. (3). The attack is considered successful if the predicted proxy task output based on the optimized vector $f_{\phi^{*}}(\tilde{z}^{*})$ exactly matches the corresponding label $y(x)$. The ASR averaged over the test dataset is $99.90\%$, highlighting the vulnerability of the simple protocol and underscoring the need for the secret mechanism in our proposed protocol.

### v2

@@ F.6 Experimental Details of Adapter Attack
F.7 Experimental Details of Secret Recovery Attack
@@
F.8 Case Study: The Vulnerability of the Simple Protocol Without Secret Mechanism
In this case study, we implement the simple protocol and examine its vulnerability to the direct vector optimization attack described in Section 3.2. We use the SoW representation as the self-labeling function. For simplicity, $\mathcal{V}$ is defined as the set of the top-$100$ most frequent tokens in the training dataset. We use Llama-2-13B as the specified model. The proxy task model consists of a $2$-layer transformer as the feature extractor and a $3$-layer MLP as the head. The model is trained for $8$ epochs with a batch size of $512$.
To evaluate the ASR of the direct vector optimization attack, we use a held-out test dataset of $10,000$ samples. Each attack vector $\tilde{z}$ is randomly initialized and optimized over $100$ steps using the Adam optimizer [20] based on Eq. (3). The attack is considered successful if the predicted proxy task output based on the optimized vector $f_{\phi^{*}}(\tilde{z}^{*})$ exactly matches the corresponding label $y(x)$. The ASR averaged over the test dataset is $99.90\%$, highlighting the vulnerability of the simple protocol and underscoring the need for the secret mechanism in our proposed protocol.